# arXiv:2603.00422v1[stat.AP] 28 Feb 2026

### 28 Feb 2026 v1

@@ 7 Conclusion
This supplement provides technical details supporting the main text. Appendix S1 describes data measurement issues. Appendix S2 gives the full simulation specification. Appendix S3 surveys the methodological building blocks for coupled forecasting. Appendix S4 formalizes the search-to-booking distance metrics.
@@
Inventory $S_{t}$ is typically constructed from calendar scrapes or platform dashboards. The primary measurement problems are: (i) distinguishing blocked dates (host-side unavailability) from booked dates (guest-side demand), since both appear as “not available” in a calendar scrape; (ii) handling minimum-stay rules, which make individual nights conditionally available depending on booking length; (iii) accounting for multi-night bookings, which occupy consecutive listing-nights but represent a single demand event; and (iv) sampling bias, since scrape frequency and coverage vary across vendors, and inactive or stealth listings inflate raw counts (Wang et al., 2024). The conceptual object $S_{t}$ is defined as inventory available at any posted price; $P_{t}$ and $R_{t}$ then filter which portion of that inventory is exposed to demand under specific search conditions.
@@
Within the Dirichlet family, Bayesian Dirichlet ARMA (BDARMA) models combine the Dirichlet likelihood with VARMA dynamics on the mean parameters after an additive log-ratio transformation (Katz et al., 2024). Extensions address time-varying volatility through a Dirichlet ARCH component (Katz and Weiss, 2025), structural breaks via a directional-shift intervention mechanism (Katz, 2026), and density forecast calibration through a centered-innovation formulation (Katz, 2025). Software implementation is available in the darma R package (Katz and Wilson, 2024).
@@
The censoring formulation $B_{t}\leq\min(D_{t},S_{t})$ connects the platform forecasting problem to a mature literature on unconstrained demand estimation in revenue management and demand censoring in inventory systems. In hotel revenue management, observed room sales are censored by available capacity, and recovering the latent demand distribution is a prerequisite for optimal pricing and allocation (Talluri and van Ryzin, 2004). The statistical problem is classical: the forecaster observes a truncated or censored outcome and must infer the parameters of the uncensored distribution. Standard approaches include expectation-maximization algorithms for censored Poisson or normal demand (Talluri and van Ryzin, 2004), booking curve detruncation methods that project partial booking curves to their uncensored completions (Schwartz, 1997; Weatherford and Bodily, 1992), and Bayesian formulations that place priors on both the demand parameters and the censoring probability.
@@
For density forecasts, Gneiting and Raftery (2007) establish the framework of proper scoring rules that reward sharpness subject to calibration. In multivariate settings, however, standard energy scores have limited sensitivity to misspecified dependence structures (Pinson and Tastu, 2013). Scheuerer and Hamill (2015) propose variogram-based proper scoring rules that are explicitly sensitive to the dependence between forecast components. For coupled supply-demand forecasting, dependence-sensitive scores are essential because joint forecast errors propagate across sides.

### 19 Apr 2026 v2

@@ 8 Conclusion
This supplement provides technical details supporting the main text. Appendix S1 expands on the measurement discussion already summarized in the main text. Appendix S2 gives the full simulation specification. Appendix S3 surveys the methodological building blocks for coupled forecasting. Appendix S4 formalizes the search-to-booking distance metrics.
@@
This appendix expands on the main-text discussion of measurement, observability, and aggregation. Inventory $S_{t}$ is typically constructed from calendar scrapes or platform dashboards. The primary measurement problems are: (i) distinguishing blocked dates (host-side unavailability) from booked dates (guest-side demand), since both appear as “not available” in a calendar scrape; (ii) handling minimum-stay rules, which make individual nights conditionally available depending on booking length; (iii) accounting for multi-night bookings, which occupy consecutive listing-nights but represent a single demand event; and (iv) sampling bias, since scrape frequency and coverage vary across vendors, and inactive or stealth listings inflate raw counts (Wang et al., 2024). The conceptual object $S_{t}$ is defined as inventory available at any posted price; $P_{t}$ and $R_{t}$ then filter which portion of that inventory is exposed to demand under specific search conditions.
@@
Within the Dirichlet family, Bayesian Dirichlet ARMA (BDARMA) models combine the Dirichlet likelihood with VARMA dynamics on the mean parameters after an additive log-ratio transformation (Katz et al., 2024). Extensions address time-varying volatility through a Dirichlet ARCH component (Katz and Weiss, 2026), structural breaks via a directional-shift intervention mechanism (Katz, 2026), and density forecast calibration through a centered-innovation formulation (Katz, 2025). Software implementation is available in the darma R package (Katz and Wilson, 2024).
@@
The censoring formulation $B_{t}\leq\min(D_{t},S_{t})$ connects the platform forecasting problem to a mature literature on unconstrained demand estimation in revenue management and demand censoring in inventory systems. In hotel revenue management, observed room sales are censored by available capacity, and recovering the latent demand distribution is a prerequisite for optimal pricing and allocation (Talluri and van Ryzin, 2004). The statistical problem is classical: the forecaster observes a truncated or censored outcome and must infer the parameters of the uncensored distribution. Standard approaches include expectation-maximization algorithms for censored Poisson or normal demand (Talluri and van Ryzin, 2004), booking curve detruncation methods that project partial booking curves to their uncensored completions (Schwartz and Hiemstra, 1997; Weatherford and Bodily, 1992), and Bayesian formulations that place priors on both the demand parameters and the censoring probability.
@@
For density forecasts, Gneiting and Raftery (2007) establish the framework of proper scoring rules that reward sharpness subject to calibration. In multivariate settings, however, standard energy scores have limited sensitivity to misspecified dependence structures Pinson and Tastu (2013). Scheuerer and Hamill (2015) propose variogram-based proper scoring rules that are explicitly sensitive to the dependence between forecast components. For coupled supply-demand forecasting, dependence-sensitive scores are essential because joint forecast errors propagate across sides.